The material point with mass m = 500 g performs rectilinear motion and is under the action of constant force F = 50 N. Calculate the change of momentum during delta t1; = 0.5 s and delta t2 = 2 s


1
Expert's answer
2021-01-28T09:56:40-0500

We can find the change of momentum from the impulse momentum change equation:


"J=\Delta p=F_{avg}\Delta t,""\Delta p=50\ N\cdot(2\ s-0.5\ s)=75\ N\cdot s."